OpenStudy (anonymous):

A muffin recipe calls for 4 tablespoons of water to make 10 servings. How many tablespoons of water are needed to make 8 servings? A. 2.5 tablespoons B. 3.2 tablespoons C. 5 tablespoons D. 20 tablespoons

OpenStudy (anonymous):

can i help yu

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

B

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

lol aajugdar

OpenStudy (anonymous):

I was also expecting a step by step

OpenStudy (anonymous):

@wanthelp According to the code of conduct, you shouldn't give direct answers

ganeshie8 (ganeshie8):

\(\large \dfrac{4}{10} = \dfrac{x}{8}\)

OpenStudy (anonymous):

hmm okay you need 4 tablespoons for 10 servings so for 1 serving you need 4/10 = 0.4 tablespoons so for 8 it will be 0.4 x 8 = ? @ganeshie8 's method is easier to do it

OpenStudy (anonymous):

3.2!

OpenStudy (anonymous):

so wanthelp was right

OpenStudy (anonymous):

B

OpenStudy (anonymous):

thank you... can i have a medal

OpenStudy (anonymous):

she was indeed lol

OpenStudy (anonymous):

lolol

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

fine....

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

my medal also @wanthelp for being so helpful :)

OpenStudy (anonymous):

thank you